OpenStudy (anonymous):

What is the equation of a circlewith the center at (2,5) and a radius of 3 units?

OpenStudy (callisto):

(x-h)^2 + (y-k)^2 =r^2 Put h=2, k=5 and r=3 to (x-h)^2 + (y-k)^2 =r^2 Can you show us what you've got?

OpenStudy (hoblos):

general form : (x-a)^2 + (y-b)^2 = r^2 where the center is (a,b) abd the radius is r

OpenStudy (anonymous):

(x-2)^2+(y-5)^2=3^2

OpenStudy (anonymous):

yea , as callisto said , please solve it for us , and if you face any prob, let us know !